### v1

# A Comprehensive Evaluation of Semantic Relation Knowledge of Pretrained Language Models and Humans

Abstract
Recently, much work has concerned itself with the enigma of what exactly PLMs (pretrained language models) learn about different aspects of language, and how they learn it. One stream of this type of research investigates the knowledge that PLMs have about semantic relations. However, many aspects of semantic relations were left unexplored. Only one relation was considered, namely hypernymy. Furthermore, previous work did not measure humans’ performance on the same task as that solved by the PLMs. This means that at this point in time, there is only an incomplete view of models’ semantic relation knowledge. To address this gap, we introduce a comprehensive evaluation framework covering five relations beyond hypernymy, namely hyponymy, holonymy, meronymy, antonymy, and synonymy. We use six metrics (two newly introduced here) for recently untreated aspects of semantic relation knowledge, namely soundness, completeness, symmetry, asymmetry, prototypicality, and distinguishability and fairly compare humans and models on the same task. Our extensive experiments involve 16 PLMs, eight masked and eight causal language models. Up to now only masked language models had been tested although causal and masked language models treat context differently. Our results reveal a significant knowledge gap between humans and models for almost all semantic relations. Antonymy is the outlier relation where all models perform reasonably well. In general, masked language models perform significantly better than causal language models. Nonetheless, both masked and causal language models are likely to confuse non-antonymy relations with antonymy.
1 Introduction
What do pretrained language models (PLMs) learn about human language? This question has recently been a central topic of discussion in Natural Language Processing (NLP) and Computational Linguistics (CL). PLMs are utilized in various situations but are not thoroughly understood. While initial work explored syntactic and factual knowledge [49, 51, 7, 32, 42], more recently there are a number of studies focusing on lexical semantic knowledge, particularly knowledge about semantic relations [16, 50, 21].
Semantic relations describe how the senses of two lexical items are related. They are an important aspect of linguistic knowledge because they structure the vocabulary of natural languages [40, 37, 55]. This makes them essential for both human language comprehension and production. On the modeling side, semantic relations are crucial for tasks such as text simplification, paraphrasing, natural language inference, and discourse analysis, as has been shown experimentally [58, 34, 20, 1]. Therefore, it is beneficial and necessary for PLMs to learn semantic relations well.
The present study tries to establish to what extent they are able to do so. We extend the existing methodology by introducing a new evaluation framework. We cover six relations, namely hypernymy, hyponymy, holonymy, meronymy, antonymy, and synonymy. Our framework is also the first to shine a light on previously understudied properties of semantic relations and meta-relations. Two kinds of comparisons are considered. First, we compare models against humans on the same task, so that we can quantify the difference with the theoretically achievable ceiling. Second, we compare two families of models, which differ in the pretraining tasks used. Within families, we consider different sizes ranging from millions to billions in parameters. These comparisons allow us to identify which factors facilitate the acquisition of semantic relation knowledge by the models. In sum, our evaluation framework deeply explores both well-studied and previously unexplored properties, and it uses new metrics in a new comparative experimental setting. By doing so, it adds both depth and width to the current knowledge about the quality of semantic relation knowledge that can be acquired by today’s PLMs.
2 Related Work
2.1 Probing for Hypernymy
Hypernymy is a typical semantic relation. In hypernymy, one word (the hyponym) refers to a specific concept and another word (the hypernym) refers to a more general concept encompassing the hyponym’s meaning. For example, “bird” is a hypernym of “robin”. Hyponomy is the name of the opposite relation: “robin” is a hyponym of “bird”.
In order to study the linguistic knowledge of PLMs, including hypernymy, several methods have been proposed in the past. Standard approaches include the use of probing classifers [24, 23, 36, 35, 4] and prompt-based probing [49, 16, 51, 7, 32].
A probing classifier is a neural classifier that takes a word embedding or a combination of word embeddings as input and determines whether a linguistic property of interest holds. In the context of probing for hypernymy knowledge, a probing classifier may take the concatenation of the embeddings of “robin” and “bird” as input and perform a binary classification, determining if hypernymy holds. The performance of the classifier can then be interpreted as the extent to which the linguistic property in question is successfully encoded in the word embeddings. Probing classifiers have the disadvantage that they require training, and that they also introduce new parameters. They can therefore encounter the problem of double interpretation, where the researcher needs to interpret two sets of parameters at the same time: one set coming from the pretrained model probed, and another from the probing classifier. Such situations lead to circularity.
In contrast, Ettinger [16] pioneered the study of hypernymy with prompt-based probing. In prompt-based probing, responses from a model are elicited using a prompt, a textual string with slots, all of which are unfilled. An example prompt for hypernymy is “a [w] is a [v] where [w] is a slot for the word given (called the target word here), whereas [v] will be predicted by a model. As the second word, which is called relatum, depends on the target word and the relation described. A probe is a prompt where the target word has been filled in. For example, Ettinger used the string “a robin is a [v] as a probe. If a PLM has learned hypernymy well, it should be able to predict hypernyms of “robin” for [v], such as “bird”.
Prompt-based probing does not introduce new parameters and so avoids the aforementioned circularity in interpretation. Moreover, since prompt-based probing is a language modeling task, it aligns well with the pretraining task of PLMs. Therefore, we consider prompt-based probing the natural choice for exploring semantic relation knowledge.
Ettinger tested two models, BERT-base and BERT-large [14]. The models’ responses were evaluated by comparison with correct answers. In Ettinger’s hypernym prediction setting, both the target word (a hyponym) and the correct answer (a hypernym) were restricted to nouns, which came from a psycholinguistic experiment conducted by Fischler et al [17]. Ettinger found that both BERT-base and BERT-large achieve an accuracy of around 0.40 and a $\mathrm{Precision@5}$ score of 1.00, but these numbers were based on only 18 target words.
Ravichander et al [50] performed similar experiments with a larger dataset, also using BERT. Some of the target words they originally wanted to use were sense-ambiguous, but they decided to remove these from the evaluation set, resulting in 576 unambiguous target words. They also measured to which degree models are affected when target words are changed from singular to plural form, and found that BERT’s accuracy dropped from 0.68 for singular target words to 0.44 for plural ones. Ravichander et al concluded from this that BERT’s hypernymy knowledge is not robust.
Hanna and Mareček [21] further developed Ettinger’s and Ravichander et al’s prediction task, exploiting the fact that a semantic relation can be expressed by multiple prompts. For the prompts, Hanna and Mareček adopted some lexico-syntactic patterns known from previous work  [22] to be effective at retrieving word pairs in hypernymy and hyponymy. Such prompts include “my favorite [w] is a [v], “a [w], such as a [v], “a [w] is a type of [v], “a [w] is a [v]. Hanna and Mareček dealt with the problem of ambiguous target words in a different way from Ravichander et al [50]. They attached an example sentence from SemCor [30] to each probe, whether the target word was ambiguous or not. The example sentence was chosen in such a way that the target word was in a specific sense, namely the first WordNet sense whose hypernyms include the gold hypernym. In Hanna and Mareček’s experiments, BERT reached an accuracy of 0.48 in the best setting. Contrary to intuition, the accuracy dropped by about 0.05 when the example sentences were used.
Another modification of the original task is the evaluation data Hanna and Mareček used, which came from category norms. Category norms are items that humans judge to be subtypes of a category given to them. For instance, one category norm for “fish” is “tuna”; “trout” and “salmon” are others. Cohen et al [11] were the first to collect category norms. In their experiment, each participant wrote down category norms included in one of 43 categories given to them, as many as came to mind in 30 seconds. Cohen et al’s original category norm collection was later expanded by Battig and Montague [3], who added 13 new categories (for a total of 56 categories and 2,082 category-category norm pairs). Another colletion effort by Overschelde et al [46] added 14 more new categories (for a total of 70 categories and 1,983 category-category norm pairs).
Most of the category norms stand in hypernymy relation with their category, but one can also find some rare cases where the category and the category norms are connected by meronymy, i.e., the part-of relationship. For example, in Overschelde et al’s data, there is a category named “part of building”, which contains the category norms “window”, “door”, and “roof”, amongst others.
Hanna and Mareček applied the hyperynym-based part of Battig and Montague’s category norm data to their probing experiments in the obvious way, using category norms as target words and categories as gold relata. They discarded categories and category norms that were tokenized into multiple tokens, ending up with a total of 863 norms paired with 25 categories.
In summary, based on the above studies we can conclude that BERT is able to predict hypernyms of target words to a reasonable degree. However, several interesting avenues have been left unresearched. The most obvious of these is that there are several other established semantic relations apart from hypernymy.
2.2 Relations beyond Hypernymy
Hypernymy is only one type of semantic relation, although it is an essential one. Other semantic relations also have been a long-standing research topic in psychological, theoretical, and computational linguistics. Studies revealed that the distinction between semantic relations is non-trivial for both humans and models [9, 10, 27, 56, 44, 2, 65].
In psycholinguistics, Chaffin and Clark [9] researched the similarities and difference between several semantic relations, as perceived by humans. The experiment used a semantic sorting task, where participants are instructed to group together 31 word pairs, each representing a particular semantic relation. The relations came from five broad categories: contrast (including antonymy), similars (including synonymy), class inclusion (including hypernymy), part-whole (holonymy), and case relations (such as the agent-instrument relation and the agent-action relations, exemplified by “farmer”/“tractor” and “dog”/“bark”). The results showed that the human subjects were able to distinguish contrast (including antonymy) most easily from the other four relations. Similars (including synonymy) and class inclusion (including hypernymy) formed a second cluster, whereas case relations and part-whole formed a separate cluster each.
The similarities between hypernymy and holonymy have been extensively discussed in the semantic literature [12, 64, 27]. Both Cruse and Winston et al pointed out that hypernymy and holonymy are similar in that they both involve division and inclusion. Word pairs related by hypernymy and holonymy always consist of a word refering to an entity that undergoes division, with the other word referring to the result of that devision, whether as a part or a subclass. Joosten considered both relations under the term denotational inclusion. The similarity between the relations becomes even more obvious when collective nouns are involved. For example, we can say that a table is a kind of furniture, and we can also say that it is a part of furniture, expressions typically associated with hypernymy and holonymy [27].
Another distinction which is well-known to be difficult is that between antonymy and other relations such as synonymy and hypernymy. For both the antonymy and synonymy relation, word pairs possess high paradigmatic similarity, i.e., the words in a pair are interchangeable. Distributional methods, which are based on co-occurrence statistics, therefore struggle with the distinction between antonymy and synonymy [41]. The problem has motivated various sophisticated technical solutions [56, 45, 18, 63].
Antonymy and hypernymy are also difficult to distinguish for unsupervised distributional measures, as was shown experimentally by Shwartz et al [57]. They used the hypernymy discrimination task, which consists of distinguishing word pairs that stand in hypernym relation, from word pairs in one other relation. Each non-hypernymy relation was tested separately by creating a mixture of word pairs in it and in hypernymy relation. The non-hypernymy relations tested are antonymy, synonymy, meronymy, and the attribute relation [^1] Attribute relation is the relation holding between an adjective and a related attribute, such as “cold” and “temperature”.. Shwartz et al compared an extensive number of unsupervised distributional measures on this task, and found that in all experiment settings, it was always the antonymy mixture that yielded the lowest performance out of all mixtures.
Synonymy and hypernymy are also closely related, as was confirmed during the creation of Hyperlex [62]. Hyperlex is a lexical resource of semantic relations (hypernymy, hyponymy, meronymy, synonymy, antonymy, co-hypernymy [^2] Co-hypernymy is the relation between two concepts that share a hypernym, such as “hawk” and “robin”.), holding between 2,616 word pairs of nouns or verbs333290 unrelated word pairs also exist.. The original pairs in Hyperlex were sampled from WordNet [39] and the University of Southern Florida Norms dataset (USF) [43]. Three human checkers were asked to verify whether the semantic relation holds for each pair sampled; the pair was discarded unless two of them agreed that it did. Next, different crowd workers were asked to assign a score to the remaining pairs, indicating the degree to which the pair satisfies hypernymy. Note that for those pairs that were related in a non-hypernymy relation, the crowd workers should assign a low score. However, the humans’ score for synonymy was close to that for hypernymy pairs if the two words were close to each other in the WordNet hierarchy (i.e., separated by at most two levels).
2.3 Symmetry, Asymmetry and Prototypicality
When assessing a model’s knowledge about semantic relations holistically, it is necessary to consider not only if the model uses the semantic relations correctly, but also to what extent it learns which specific properties the semantic relations have.
Symmetry and Asymmetry
One such property is symmetry, which is defined as follows: If a word pair $(w_{1},w_{2})$ is in a symmetric relation, then the reverse pair $(w_{2},w_{1})$ is also in the relation. A related property is asymmetry, the opposite of symmetry: if the word pair $(w_{1},w_{2})$ is in an asymmetric relation, then the reverse pair $(w_{2},w_{1})$ is not in the relation. It has been experimentally shown that modeling asymmetry improves the performance of semantic relation classification tasks [19].
Symmetry and asymmery are also properties of some factual knowledge relations. For instance, the factual knowledge relation “is a sibling of” is symmetric. Mruthyunjaya et al [42] proposed metrics in order to assess whether models learn properties of such factual relations, including symmetry. They used prompt-based probing and defined the concept of reciprocal elicitation: for any word pair (target word $w$ and relatum $v$) that forms a symmetric relation, the model should respond with the relatum, when given the target word, and also respond with the target word, when given the relatum. For the probe “Bart Simpson is a sibling of [v], they expected models to predict “Lisa Simpson”, and for the converse probe “Lisa Simpson is a sibling of [v], to predict “Bart Simpson” .
For every such word pair $(w,v)$, a prompt $p$, and a model $m$’s top $k$ items in the response $m_{k}(w,p)$, given $w$ and $p$, the symmetry score is the average of $\mu_{k}(w,v,p)$ over all $(w,v)$ and $p$.
$\mu_{k}(w,v,p;m)=\mathbb{I}[v\in m_{k}(w,p)]\times\mathbb{I}[w\in m_{k}(v,p)]$ (1)
where $\mathbb{I}[P]$ is the indicator function that becomes one when $P$ is true and zero otherwise. In contrast, for asymmetric relations there should be no reciprocal elicitation but only forward elicitation: the model should only respond with the relatum given the target word, not with the target word given the relatum. The asymmetry score is the average of $\alpha(w,v,p)$ over all $(w,v)$ and $p$.
$\alpha_{k}(w,v,p;m)=\mathbb{I}[v\in m_{k}(w,p)]\times\mathbb{I}[w\notin m_{k}(v,p)]$ (2)
Both scores range between zero and one. Intuitively, the symmetry score can be interpreted as the probability of reciprocal elicitation for a semantic relation $r$. The asymmetry score, on the other hand, can be interpreted as the probability of only forward elicitation happening for $r$. For symmetric relations, a high symmetry score means that agents were able to detect symmetry; for asymmetric relations, a high asymmetry score similarly means that agents were able to recognize the asymmetric nature of the relation.
Mruthyunjaya et al’s results showed that, for symmetry, BERT outperforms even GPT-3 [6]. For asymmetry, GPT-3 outperforms BERT.
Prototypicality
Another property of interest is prototypicality [54, 52, 53]. Rosch [52] posited that not all members of a category are equally exemplary of the category, but that there is a prototype, which is the best examplar among the members. The prototypicality of any member of a category is then the degree to which it is exemplary of its category. Rosch [53] empirically found that, among the category norms of “bird” established by Battig and Montague, “robin” is the prototype, that “penguin” has the lowest prototypicality, and that “raven” and “parrot” are somewhere in the middle. There is a close relationship between prototypicality and hypernymy/hyponymy, as is implicit in the construction of her experiment [^4] Despite this obvious relationship, Rosch did not explicitly use the term hypernymy..
There have been theoretical discussions of the prototypicality of holonymy. Taylor [59] and Tversky [61] presented top-down accounts, which emphasize that the whole (holonym) is intrinsic in the conceptualization of the part (meronym). They therefore predict that the relationship between the whole and its mandatory parts should be tighter than between the whole and its optional parts. A building is a structure with walls, and walls are defined by their function within a building. This makes “building” a prototypical holonym of “wall”. But not all holonymy pairs are well described by these accounts, because some optional parts also play an important role. This is acknowledged in the bottom-up accounts [31, 38, both cited by Joosten [27]], who state that a whole is formed by assembling a number of other individual wholes, each of which has a separate existence outside the holonymy relation. For example, “sky” is a typical holonym of “cloud” (indeed, in our forthcoming experiments, it happens to be the holonym most frequently mentioned by humans). However, in sunny weather, the sky can be cloudless, so clouds are not intrinsic to the sky. This seems to make bottom-up theories descriptively more adequate, although neither proposes a prototype prediction mechanism.
The antonymy relation also shows prototypicality effects; this has been empirically confirmed with corpus experiments [26] and human experiments [47, 48]. The two other semantic relations of interest to us (meronymy and synonymy) have not been studied in connection with prototypicality, either theoretically and empirically. There are also no experimental studies that evaluate how neural models learn prototypicality for any relation.
3 Methodological Considerations
The majority of the previous research studied hypernymy. Beyond hypernymy, there is also much theoretical and experimental knowledge about semantic relations. Additionally, the literature has established several facts about meta-relations. However, when it comes to practical investigations of model behaviour, it is always only hypernymy that has been studied, even though it is merely one semantic relation amongst many.
Additionally, we have seen that the relevant theoretical literature has extensively studied meta-relations such as confusability between semantic relations (cf. section 2.2). In stark contrast with this, the methodology previously used is unable to establish the degree to which a model mistakes one semantic relation for another, and methodologies for other meta-relations are non-existent. This leaves us with an incomplete understanding of the nature of semantic relations, and of the knowledge that PLMs have about semantic relations.
Therefore, we design new metrics for measuring prototypicality (a property of relations) and distinguishability (a meta-relation). Using these new metrics, and the established ones for symmetry and asymmetry, we study hyponymy, holonymy, meronymy, antonymy and synonymy, as well as hypernymy. We further provide a direct comparison between models and humans on the same task. Such a human ceiling will allow us to interpret the performance of models more meaningfully.
Word Senses
Another recurrent problem for all probing experiments is that most target words from any source, are naturally sense-ambiguous words. Previous experimental attempts to deal with sense ambiguity are suboptimal. Ravichander et al limited the target words they use to the unambiguous words. This limits the research focus to an artificial subset of all possible words and relations, and has the practical disadvantage of considerably reducing the number of target words one can use. The other existing solution is to provide context of the target word sense, for example in the form of example sentences, as Hanna and Mareček did, but this empirically harmed the performance of models. Our methodology offers an alternative solution to this problem.
Introduction of Relatum Sets
Some previous experiments on relatum prediction assumed that there is only one correct relatum for each tuple. Hanna and Mareček [21] acknowledge that “orange” has two gold relata: “color” and “fruit”. They therefore defined two separate gold tuples for this probe: (“orange”, HYP, “color”) and (“orange”, HYP, “fruit”). However, when calculating the accuracy scores, they consider only the first item in the responses for both tuples, and then average over the two tuples. In this setting, it is theoretically impossible for a model to achieve the full score (1) for any sense-ambiguous probe, even if the model had the ability to predict both relata. Whether the model predicts [“color”, “fruit”] or [“fruit”, “color”], the accuracy score is always 0.5555In the general case, the highest achievable accuracy score for target words with $n$ relata is $\frac{1}{n}$..
This means that the extent to which a model can predict all relata of a target word is undervalued. The gold standard we want to define should treat target words with multiple relata more fairly. We define gold standards as a set of relata, which we call relatum set. Working with relatum sets is particularly necessary when evaluating hyponymy and meronymy, since in these two relations multiple relata cases are likely to particularly frequent. Under the use of multiple relata, accuracy alone is no longer suitable for evaluation; instead, metrics borrowed from information retrieval are required.
The introduction of relatum sets has an important side-effect in that it enables the evaluation of models’ recognition of prototypicality. However, we do not know a priori if prototypicality holds for all relations of interest. This question needs to be experimentally established. Once the prototypicality of a relation is confirmed, the relatum set allows us to measure the degree to which the model has captured prototypicality, by a comparison to the human responses.
Determiners
A confounder in the interpretation of numerical results is the use of definite and indefinite determiners in the probes. Previous researchers routinely used probes that include indefinite determiners, such as Ettinger’s (2020) “a robin is a [v]. The English indefinite determiner changes its form from “a” to “an” if the following word’s pronunciation starts with a vowel. Choosing “a” or “an” in a probe before the [v] slot would therefore bias the prediction by models towards relata with an initial vowel or consonant.
Ettinger ran experiments using both types of determiners, comparing pairs of probes differing only in the determiners used. Manual inspection of the responses showed that BERT indeed always adhered to the morphophonetic rule. For example, given the probe “a hammer is a [v], BERT-large predicts [“hammer”, “tool”, “weapon”, “nail”, “device”], whereas given the probe “a hammer is an [v], it predicts [“object”, “instrument”, “axe”, “implement”, “explosive”]. This suggests that the BERT models were able to utilize the grammatical information contained in the determiner as a clue.
Ravichander et al, when faced the problem of which indefinite determiner to place before [v], chose to always use the determiner that morphophonetically fits with the gold standard answer. For example, the determiner in the probe “a moth is an [v] was chosen to be “an” exactly because the gold standard answer was “insect”. However, when the determiner acts as a clue for the model, it becomes impossible to disentangle how much of the results is due to the model’s semantic knowledge and how much to the clue. Choosing randomly also doesn’t solve the problem. Our solution uses both probes with “a” and probes with “an” and merges the results in a statistical manner.
Model Comparison
Another understudied aspect is the comparison between different types of PLM models on the semantic probing task. BERT, the only model studied so far, is a masked language model (MLM). MLMs are pretrained on the masked language modeling task, in which a model is asked to recover tokens in a given sentence that are randomly masked. But recently, causal language models (CLMs) such as OPT [66] and Llama [60] have shown high performance in many tasks and thus gained attention. CLMs are pretrained on next-token prediction, the task of predicting the rightmost word given a sequence of words. The difference in pretraining tasks means that the models make their decisions based on different kinds of information. MLMs consider the context of both sides of the masked word, while CLMs consider only the preceding context. Previous studies in a number of tasks found a large influence of the type of context used in PLM pretraining on performance. For instance, for factual knowledge, MLMs have been found to be superior over CLMs [49, 8, 42]. We are the first to study the role of bidirectional contexts in the recognition of semantic relations. To be fair to both MLMs and CLMs, we have designed all our prompts in such a way that they end with the slot [v].
Model Size
Apart from model type, model size may also matter. For pretraining tasks, Kaplan et al [28] showed that if the corpus size is fixed, larger models show smaller losses and thus better performance. This regularity was found to be empirically valid for other sentence completion tasks [6], not only for pretraining tasks. However, there are contrary reports from factual relation recognition that smaller models (BERT and RoBERTa) outperform larger models (GPT-4 and GPT-3) in the determination of some properties [42]. When it comes to semantic relation tasks in general, it is unknown which of these tendencies is stronger.
We always first establish human performance for each task and then use it as the measuring stick for models’ performance [^6] Except in the case of asymmetry, as we will explain in what follows..
The rest of the article is structured as follows. Section 4 describes the material collection. We will then describe our proposed evaluation metrics in Section 5. The following Sections 6 and 7 will present the settings of human and model experiments, with results following in Section 8.
4 Data
Our evaluation employs prompt-based probing. In order to carry out the evaluation, we need prompts, target words and a gold-standard relatum set for each target word. We will now explain how we collected them.
4.1 Prompt Design
The underlying object we operate over is called a word-relation-relatum tuple (tuple in short). We denote it by $t^{r}=(w,r,v)$, where $r\in R$ is a semantic relation, $w$ is a target word and $v$ is an $r$-relatum, i.e., a word standing in relation $r$ to $w$ [^7] In what follows, we omit the $r$- in the term if it is clear which specific relation is meant.. $T^{r}$ is the set of such tuples. Our set of relations $R$ consists of hypernymy (HYP), hyponymy (HPO), holonymy (HOL), meronymy (MER), antonymy (ANT), and synonymy (SYN).
For each relation $r$, we construct a set of prompts $p^{r}\in P^{r}$. We reuse Hanna and Mareček’s prompts for hypernymy. For the other relations, we hand-craft new prompts. In total, we use seven prompts for hypernymy, synonymy and holonymy; four prompts for hyponymy; six prompts for meronymy; and nine prompts for antonymy [^8] A full list of prompts can be found in Appendix A..
Examples follow.
[…]
Note that some prompts do not require any determiner, but others do. The notation [det] expresses that either “an” or “a” is chosen, based on certain conditions to be discussed later. Our prompts are formulated such that [w] always preceeds [v], and that there is no token after the [v] slot.
4.2 Target Words and Probes
Tuples
In order to obtain our set of tuples $T^{r}$, we use existing word-relation-relatum tuples from Hyperlex [62] and from the category norm corpus by Overschelde et al [46]. We use only those tuples where both $w$ and $v$ are nouns, and where both are contained in the intersection of the vocabularies of all models that will be tested in the experiment. This resulted in a total of 1,347 tuples: 147 for hypernymy, 805 for hyponymy, 234 for meronymy, 52 for antonymy and 109 for synonymy. Note that none of the sources contributed any holonymy tuples.
To get more tuples, we expand our set of tuples by symmetric augmentation. Symmetric augmentation can be applied to symmetric relations (here: antonymy and synonymy) by adding tuples where $v$ and $w$ are swapped, as follows:
$T^{r,aug}=T^{r}\cup\{(w,r,v)\mid\forall(v,r,w)\in T^{r}\}.$ (4)
Symmetric augmentation can also be applied to those relations that have a reverse relation. If $(r_{1},r_{2})$ is a pair of reverse relations, the following holds:
$(w,r_{1},v)\in T^{r_{1}}\iff(v,r_{2},w)\in T^{r_{2}}.$ (5)
Hypernymy and hyponymy form a reverse relation pair; holonymy and meronymy form another reverse relation pair. A small change to the symmetric augmentation procedure is necessary. For relations $r_{1},r_{2}$ in a reverse relation pair, symmetric augmentation proceeds as follows:
$T^{{r_{1}},aug}=T^{r_{1}}\cup\{(v,r_{1},w)\mid\forall(w,r_{2},v)\in T^{r_{2}}\},$ (6)
$T^{{r_{2}},aug}=T^{r_{2}}\cup\{(v,r_{2},w)\mid\forall(w,r_{1},v)\in T^{r_{1}}\}.$ (7)
For example, we can reverse the meronymy tuple (“building”, MER, “wall”) to obtain a new holonymy tuple (“wall”, HOL, “building”). Note that if there are any duplicate tuples, they are removed to form the set $T^{r,aug}$. We will simplify notation after augmentation and use $T^{r}$ to refer to $T^{{r},aug}$. After symmetric augmentation, the total number of tuples has risen 1.66 fold (2,242, from 1,347). For holonomy, this process creates the only tuples in existence (186 tuples).
Target words
Target words can now be extracted from tuples in the obvious way. For each relation $r$, we form $W^{r}$ as the set of target words $w$ from all tuples in $T^{r}$. Except for duplicates, each tuple contributes a target word. $W$, the union of $W^{r}$ for different relations $r$, denotes the set of unique target words in all experiments, independent of relation.
Probes
A probe $x^{r}\in X^{r}=\{\nu(w^{r},p^{r})\mid\forall w^{r}\in W^{r},\forall p^{r}\in P^{r}\}$ is then a string created by applying the verbalization function $\nu$ to a target word $w^{r}$ and a prompt $p^{r}$. The verbalization function $\nu$ always assigns $w^{r}$ to the first slot [w] and leaves the second slot [v] empty. An example is
[…]
For the determiner before the target word, the function selects the morphophonetically correct form, as is uncontroversial and commonly done in previous work. For the indefinite determiner before [v], more thought is required. We explain our treatment in Section 7.2.
The total number of probes we create is 10,546; for each relation, the component is the product of prompts and target words. Statistics for each relation can be gleaned from Table 1999The total over target words reported in the table is the sum over $|W^{r}|$. Because some target words are associated with more than one relation, this sum is different from $W$, the total number of unique target words, which is 1,266..
Table 1: Statistics of prompts, target words, and probes after augmentation.
Relation $|P^{r}|$ (prompts) $|W^{r}|$ (target words) $|X^{r}|$ (probes)
Hypernymy (HYP) 7 692 4,844
Hyponymy (HPO) 4 310 1,240
Holonymy (HOL) 7 186 1,302
Meronymy (MER) 6 144 864
Antonymy (ANT) 9 91 819
Synonymy (SYN) 7 211 1,477
TOTAL 40 1,634 10,546
4.3 Relatum Sets
So far, we have constructed probes that we will give as inputs to models and as stimuli to humans. We now want to create $r$-relatum sets $Y^{r}$ that we can use for evaluation, for each target word $w\in W$ and relation $r$. We start by considering which properties good gold standards for our task would have.
First, we want a sufficient number of $r$-relata for each target word in $W$. This is important for a fair evaluation of relatum prediction ability. If a dataset has only few relata per target word, we are lacking information about what the potential relatum set could look like, resulting in sparse data bias. We therefore need larger relatum sets.
Second, each target word should be associated with as many relations as possible. One of the abilities that we are going to evaluate is the degree to which models and humans can distinguish relations from each other. In principle, the more relations are present, the better the resulting evaluation should be. At a minimum, each target word needs to be associated with two relations to make this measurement possible; at a maximum, each relation can be confused with five other relations.
Unfortunately, the current relatum sets do not fulfill these two criteria. On average, they contain only 1.2 (antonymy) to 2.6 (hyponymy) relata per target word, and the average number of relations associated with each target word is only 1.3. This means that a majority of target words is associated with only one relation, making it impossible to assess models’ ability to distinguish between relations. Therefore, it is desirable to increase the number of relata per relatum set, as well as the number of associated relations for each target word.
For a given relation $r$ and a target word $w$, we increase the $r$-relata in the $r$-relatum set $Y^{r}$ as follows. We first retrieve all possible word senses of $w$ in WordNet. Then, for each word sense that has $r$-relata documented in WordNet, we update $Y^{r}$ by adding the new $r$-relatum, unless it is not included in the models’ shared vocabulary. In other words, all possible $r$-relata of any sense of the target word are included in the expanded $Y^{r}$. For example, “ending/1” is a synonym of “termination/4”, and “ending/3” is a synonym of “conclusion/6”. The resulting synonym set for ending therefore includes both “termination” and “conclusion”. This is so despite the fact that they refer to different senses of “ending” [^10] Note that the construction of synonym set for “termination” or “conclusion” results in a different expanded synonym set.. We additionally include indirect hypernyms and hyponyms of either of the target words senses, defined as those which lie within a path length of two in the WordNet hierarchy.
This procedure can result in a situation where more than one semantic relation holds between two word forms. For example, WordNet lists “conclusion/3” as a hyponym of “ending/3” and at the same time lists “conclusion/4” as a synonym of “ending/4”. The result are two relations holding between the word forms “conclusion” and “ending” [^11] The problem arises because our evaluation is performed at the word form level (as is the common approach), and not the sense level. If we were able to evaluate with senses disambiguated, we would be able to leave these relata in, with added profit.. We therefore solve this problem by removing all relationally ambiguous relata for each target word. After this step, we have a guarantee that for each target word, the relatum sets of different relations are mutually exclusive.
Table 2: Sizes of relatum sets before and after expansion per relation.
Relation Before Expansion After Expansion
Hypernymy (HYP) 1.2 $\pm$ 0.5 8.4 $\pm$ 6.8
Hyponymy (HPO) 2.6 $\pm$ 4.0 39.2 $\pm$ 54.9
Holonymy (HOL) 1.3 $\pm$ 0.6 2.9 $\pm$ 2.3
Meronymy (MER) 1.7 $\pm$ 3.4 3.9 $\pm$ 6.2
Antonymy (ANT) 1.1 $\pm$ 0.3 1.2 $\pm$ 0.5
Synonymy (SYN) 1.1 $\pm$ 0.2 3.5 $\pm$ 2.8
The expansion results in an increase in the average relatum set sizes, as can be seen from Table 2. The final average relatum set sizes range from 1.2 for antonymy to 39.2 for hyponymy. The average size of expanded hyponym sets is far larger than others because of the nature of hyponymy; as we descend the WordNet hierarchy to retrieve hyponyms, the number of hyponyms increases. The expansion also increases the average number of relations associated with each target word from 1.3 to 3.4.
These relatum sets constitute our gold standard in the upcoming evaluation. Of course it is possible for both humans and models to respond with a word that is not in the $r$-relatum set for either relation $r$. We call such words OOR (out of relatum set).
5 Metrics
The proposed evaluation framework consists of five metrics, two of which are novel. The novel metrics are called prototypicality and distinguishability. Prototypicality evaluates a property of semantic relations. Distinguishability evaluates agents’ ability to distinguish relations from each other. Soundness and completeness measure the performance of relatum prediction under the multiple relata setting. Symmetry and asymmetry have been studied before, but only with factual relations.
To calculate all metrics, we need a ranked list, for humans and for each model. The process starts with a probe, which we gain from target word $w^{r}$ and prompt $p^{r}$ by the verbalization function $\nu$. Using the probe, we elicit relata $v$ from multiple human participants, or from each model.
We treat the group of humans and each model as a random agent $m$. During probing experiments, models naturally produce a distribution $D(w^{r},p^{r};m)$, where each vocabulary item is associated with a probability estimate. We transform the relata coming from multiple human participants into a single comparable distribution. We do this by calculating the normalized frequency over relata, after pooling the data coming from different participants.
From each $D(w^{r},p^{r};m)$ for either agent $m$, we can create a ranked list $L(w^{r},p^{r};m)$. The rank is established by the probability of that word from our distribution over relata $D(w^{r},p^{r};m)$. The list returned by models is as long as their vocabulary, so we need to introduce a cutoff $k$, which will be established separately for each metric. We denote $L_{k}(w^{r},p^{r};m)$ as the resulting response list for either agent $m$ is denoted. This allows us to treat human responses and model responses in a comparable way.
5.1 Soundness and Completeness
Soundness and completeness are akin to precision and recall. Soundness, denoted by $\mathcal{S}(r;m)$, is the extent to which words predicted by $m$ are valid relata for relation $r$.
$\mathcal{S}(w^{r};m)$ $=\frac{1}{\lvert P^{r}\rvert}\sum_{p^{r}\in P^{r}}\mathrm{Precision@1}\big{(}L(w^{r},p^{r};m),Y^{r}\big{)},$ (11)
$\mathcal{S}(r;m)$ $=\frac{1}{\lvert W^{r}\rvert}\sum_{w^{r}\in W^{r}}\mathcal{S}(w^{r};m).$ (12)
Note that we first average $\mathrm{Precision@1}$ scores over different prompts for the same target word. We then average over target words. Others before us [16, 50, 21] have used accuracy, which is mathematically identical to $\mathrm{Precision@1}$, but their numerical values are not comparable to our soundness. This is because we use relatum sets instead of single gold standard items. As any item of our relatum set counts as a hit, soundness values will be generally higher than the accuracy values in previous work.
Completeness $\mathcal{C}(r;m)$ measures the extent to which $m$ can predict all relata for relation $r$.
$\mathcal{C}(w^{r};m)$ $=\frac{1}{\lvert P^{r}\rvert}\sum_{p^{r}\in P^{r}}\mathrm{Recall@k}\big{(}L(w^{r},p^{r};m),Y^{r}\big{)},$ (13)
$\mathcal{C}(r;m)$ $=\frac{1}{\lvert W^{r}\rvert}\sum_{w^{r}\in W^{r}}\mathcal{C}(w^{r};m).$ (14)
Completeness averages over the well-known information retrieval metric $\mathrm{Recall@k}$. Here, we set $k$ to the size of the relatum set or the size of the response list, whichever is smaller. Soundness and completeness values become identical when a relatum set only has one relatum.
5.2 Symmetry and Asymmetry
We measure whether reciprocal elicitation and forward elicitation happens for triplet $(w,r,v)$ using metrics proposed by Mruthyunjaya et al [42] (Equations (1) and (2), respectively). Our metrics differ from Mruthyunjaya et al’s in how averaging takes place. We average as we do in our calculation of the $\mathcal{S}$ and $\mathcal{C}$ scores. The summary statistic symmetry $\mathcal{M}_{k}(r;m)$ is achieved by first obtaining a symmetry score for each triplet, agent $m$, and relation $r$, and then averaging over the triplets.
$\mathcal{M}_{k}(w,r,v;m)$ $=\frac{1}{\lvert P^{r}\rvert}\sum_{p^{r}\in P^{r}}\mu_{k}(w,v,p^{r};m)$ (15)
$\mathcal{M}_{k}(r;m)$ $=\frac{1}{\lvert T^{r}\rvert}\sum_{(w,r,v)\in T^{r}}\mathcal{M}_{k}(w,r,v;m)$ (16)
where $r$ is either antonymy or synonymy and $\mathcal{M}_{k}(w,r,v;m)$ is the symmetry score for triplet $(w,r,v)$.
Asymmetry $\mathcal{A}_{k}(r;m)$ of agent $m$ on relation $r$ is calculated similarly as $\mathcal{M}_{k}(r;m)$.
$\mathcal{A}_{k}(w,r,v;m)$ $=\frac{1}{\lvert P^{r}\rvert}\sum_{p^{r}\in P^{r}}\alpha_{k}(w,v,p^{r};m)$ (17)
$\mathcal{A}_{k}(r;m)$ $=\frac{1}{\lvert T^{r}\rvert}\sum_{(w,r,v)\in T^{r}}\mathcal{A}_{k}(w,r,v;m)$ (18)
where $r$ is either hypernymy, hyponymy, holonymy, or meronymy.
For symmetric relations, the data at hand allows us to directly measure symmetry scores, for both models and humans. For asymmetric relations, we want to know if models successfully show only forward elicitation, and no backward elicitation on the same tuple. Note that for this we need a new probe, a trick probe, which expresses the opposite relationship to the original probe: on the same prompt template, it presents the relatum and then records whether backward elicitation happens, which it shouldn’t. For example, for the holonymy tuple $t^{\mathrm{HOL}}=$(“wall”, HOL, “building”), we need to create the trick probe “a building is a part of [det] [v] (in addition to what we already have, namely “a wall is a part of [det] [v]).
Trick probes can be constructed on the fly: the prompt from the original probe is used, and the target word is replaced with the relatum. This procedure yields 310$\times$7 trick probes for hypernymy, 692$\times$4 for hyponymy, 144$\times$7 for holonymy, and 186$\times$6 for meronymy.
On asymmetry, we evaluate only models. We assume that it is not necessary to evaluate humans, as they should have strong enough intuitions about the asymmetry of asymmetric relations.
5.3 Prototypicality
We now introduce metrics in order to determine the degree to which prototypicality is observed in the human responses. Note that unlike in our other experiments, we obtain human experimental performance that establishes the gold standard, rather than some pre-existing lexical data.
5.3.1 Response Entropy
Recall from section 2 that prototypicality is defined as the extent to which a particular relata is more exemplary than others, given a relation and a target word. We can quantify this in the form of the normalized entropy $\mathcal{R}(w^{r},p^{r})$ of $D(w^{r},p^{r};h)$, the distribution over relata for target word $w^{r}$ and prompt $p^{r}$ produced by humans ($h$) [^12] Normalization relies on the fact that $\log_{2}\lvert D(w^{r},p^{r};h)\rvert$ is the maximum entropy of a categorical distribution taking $\lvert D(w^{r},p^{r};h)\rvert$ categories..
$\mathcal{R}(w^{r},p^{r})=\begin{cases}-\sum_{v\in D(w^{r},p^{r};h)}\Pr(v)\frac{\log_{2}\Pr(v)}{\log_{2}\lvert D(w^{r},p^{r};h)\rvert},&\text{otherwise}\\ 0.&\text{if }\lvert D(w^{r},p^{r};h)\rvert=1\end{cases}$ (19)
As for all entropy-based metrics, lower numbers correspond to a stronger prototypicality effect. Maximal response entropy corresponds to a situation where all participants reply with the same single word, and nothing else. We define $\mathcal{R}(w^{r},p^{r})=0$ for this case [^13] The formula cannot be used in this case, as $\lvert D(w^{r},p^{r};h)\rvert=1$ and therefore $\log_{2}\lvert D(w^{r},p^{r};h)\rvert=0$.. Therefore, $\mathcal{R}(w^{r},p^{r})$ has a range between zero and one.
5.3.2 Prototypicality Score
We evaluate prototypicality of the model response by comparing the response with the human gold standard from above. In order to realize this, we need a similarity score that rewards models for satisfying the following requirements: 1) the prototype of the human response, i.e., the word most frequently elicited, is ranked highest in the model’s response, and 2) additionally, in the model’s response there are as many other words elicited from humans as possible, with similar rankings.
For example, consider the hypernymy probe “a wall is a part of [det] [v]. The human response is [“building”, “home”, “house”, “room”, …]. Given this ranked list, any model that returns the prototypical holonym “building” at the top position fulfills the first requirement. Concerning the other requirement, [“room”, “building”, “home”, “house”] is preferable to [“room”, “house”, “home”, “building”] because it preserves the precedence of “building” over “home” and “house” in the human response.
The first requirement can be implemented using the indicator function $\mathbb{I}[P]$. The second requirement can be implemented with the edit similarity $E(a,b)$ between word sequences $a$ and $b$, which is based on the edit distance [^14] The version of edit distance that we use here operates with insertion, deletion and substitution, with a weight of two for substitution and a weight of one for the other operations. The maximum value of edit distance is the weight of substitution times the length of the larger of the two sequences compared, which is bounded by $2k$. We therefore normalize the score by $2k$ and turn it into a similarity metric by reporting the distance from 1.. This time, $k$ differs for each probe; it is set to the number of words in human response for the probe in evaluation.
We thus define prototypicality $\mathcal{P}(r;m)$ as a distance metric as follows.
$\rho(w^{r},p^{r};m)$ $=\frac{1}{2}\,\mathbb{I}\big{[}L_{1}(w^{r},p^{r};m)=L_{1}(w^{r},p^{r};h)\big{]}$ (20)
$\quad+\frac{1}{2}\,E\big{(}L_{k}(w^{r},p^{r};m),L_{k}(w^{r},p^{r};h)\big{)},$ (21)
$\mathcal{P}(w^{r};m)$ $=\frac{1}{\lvert P^{r}\rvert}\sum_{p^{r}\in P^{r}}\rho(w^{r},p^{r};m),$ (22)
$\mathcal{P}(r;m)$ $=\frac{1}{\lvert W^{r}\rvert}\sum_{w^{r}\in W^{r}}\mathcal{P}(w^{r};m).$ (23)
The resulting prototypicality metric ranges between zero and one. A higher value means that a model’s response more closely resembles the human response, with a score of one meaning that it is identical to the human response.
5.4 Distinguishability
If a model distinguishes relation $r$ well from relation $s$, then the ranks of $r$-relata in the response should be overall much lower than the ranks of $s$-relata. This takes into account an aspect that soundness does not. Consider the example in Figure 1. There are two responses, A and B, to the holonymy probe “A wall is a part of [det] [v].
Figure 1: Disinguishability example for a holonymy probe.
Holonyms of “wall” are shown in green. Note that both responses have correctly placed a holonym in the top rank. Despite this, A is intuitively a better response than B because both holonyms in A are ranked before all incorrect relata, such as “arch” (a meronym) and “partition” (a hypernym). In contrast, the agent who produced B was less able to distinguish holonymy from meronymy and hypernymy. Distinguishability was designed to detect the difference between A and B.
Distinguishability Score
For an ordered pair of semantic relations $(r,s)$, we define $\delta(w^{r},p^{r},s;m)$ as mean relative rank of $s$-relata in the response to a probe of relation $r$. In addition, $\delta(r,s;m)$ is defined as the average of $\delta(w^{r},p^{r},s;m)$ over all prompts and target words. The distinguishability of $r$ from $s$, denoted by $\mathcal{D}(r,s;m)$, is the difference between $\delta(r,s;m)$ and $\delta(r,r;m)$ as follows.
$\delta(w^{r},p^{r},s;m)$ $=\frac{1}{|Y^{s}|}\sum_{v\in Y^{s}}{\rho\big{(}v,L_{k}(w^{r},p^{r};m)\big{)},}$ (24)
$\delta(r,s;m)$ $=\frac{1}{\lvert W^{r}\rvert}\frac{1}{\lvert P^{r}\rvert}\sum_{w^{r}\in W^{r}}\sum_{p^{r}\in P^{r}}\delta(w^{r},p^{r},s;m),$ (25)
$\mathcal{D}(r,s;m)$ $=\max\Big{(}\delta(r,s;m)-{\delta(r,r;m)},0\Big{)},$ (26)
where $\rho(a,b)$ is the normalized relative rank of a word $a$ in a list $b$. Normalization of $\rho(a,b)$ by $k$ (here set to the size of the relatum set of target word $w$) results in a range $[0,1]$.
Higher $\mathcal{D}$ scores indicate better distinction. Note that this metric can incur negative values, namely if the highest-ranked correct relatum is ranked after an incorrect relatum. In that case, the model has commited an error so grave that we are no longer interested in the rest of the response. We therefore assign zero to all cases of negative difference.
Note that our earlier process ensured that all relatum sets are mutually exclusive. If the intersection between the relatum sets of two relations $r$ and $s$ was not empty, any intersection item would wrongly contribute to both $\delta(r,s;m)$ and $\delta(r,r;m)$. This leads to a deflation of $\delta(r,s;m)$, meaning that the theoretically highest distinguishability cannot be reached even if an agent were able to perfectly separate $r$-relata from $s$-relata. The higher the intersection item is ranked, the stronger the negative effect becomes.
Area under the Distinguishability Curve (AuDC)
We define the area under the distinguishability curve as a summary statistic for distinguishability. The distinguishability curve is created as follows.
$\eta(p;m)$ $=\sum_{(r,s)\in R\times R\setminus{\{r\}}}{{\mathbb{I}\Big{[}\mathcal{D}(r,s;m)>p\Big{]}}}$ (27)
$\eta(p;m)$ is the number of relation pairs in $R\times R\setminus{\{r\}}$ whose $\mathcal{D}$ score is greater than a threshold $p$. $p$ can be read as the point at which we are satisfied that agent $m$ successfully distinguishes two relations, with a higher $\eta(p;m)$ requiring better distinguishability. The distinguishability curve then visualizes the relationship between $p$ and $\eta(p;m)$.
When $p=0$, all relation pairs with a positive $\mathcal{D}$ score contribute to the $\eta(0;m)$, making it maximal. As $p$ increases, fewer relation pairs contribute, resulting in a monotonic decrease in $\eta(p;m)$. At $p=1$, no relation pairs remain and the curve converges to zero. Note the similarities to the precision-recall curve in information retrieval, which is also established by varying a threshold.
The area under the distinguishability curve is obtained as follows.
[…]
AuDC ranges from zero to the number of all relation pairs, which is 30 in our case. In contrast to $\eta(p;m)$, which reflects the number of distinguishable relation pairs given a specific $p$, it reflects how many pairs an agent can distinguish on average, with higher numbers meaning higher distinguishability ability.
6 Human Experiment
We now move to the experiments, starting with the human probing experiment.
6.1 Elicitation of Human Responses
In order to collect responses to probes from human participants, we use the Amazon Mechanical Turk (MTurk) crowdsourcing platform. Participants were restricted to those 1) who have the MTurk Master qualification and currently live in either the United States, the United Kingdom, Australia, or Canada, and 2) additionally whose answers are approved more than 500 times at an approval rate above 95%. In total, 48 qualified participants were recruited.
We split the 10,546 probes from Table 1 into 276 subsets of 38 probes on average, making sure that no subset contained more than one probe with the same relation and the same target word. The time limit for responding to each probe was three minutes. Four participants were assigned to each subset. Participants answered 22 subsets on average.
We asked participants to type up to five relata for each probe. We instructed them that they should use nouns, but no multi-word expressions. We further told participants that the relata could start with either a consonant or a vowel. In addition to the real probes, we used three additional bogus probes (such as “The earth rotates around the [v]) per subset, and rejected subsets where the bogus item was not answered correctly (in this case, only “sun” was accepted). All participants correctly answered all bogus probes, so we were able to accept all responses.
The humans responded with 93,120 word tokens (7,216 word types) in total, of which 10,895 word tokens (18%) are OOR (3,691 OOR types, 51%). 2,106 response lists consist solely of OOR words (20%). On average, we find the first non-OOR words at rank 1.5 in a response list.
On the basis of all responses including OOR, we calculate soundness, completeness, symmetry for antonymy and synonymy (but not for asymmetry), prototypicality, and distinguishability. Our metrics will penalize agents for responding with an OOR word in each case.
6.2 Response Entropy Analysis
To remind the reader, our gold standard for prototypicality, unlike that for the other metrics, is an outcome of the human experiments, so needs to be calculated before model evaluation can take place.
We first give some examples of the kinds of prototypes the participants produced. For the target word “wall” under holonymy, “building” is the most prototypical item in the response, and for “cloud”, it is “sky”. These tendencies hold irrespective of which prompt was used. These two pairs reflect the top-down and bottom-up accounts of holonymy prototype theories (cf. Section 2).
We then look at hypernymy. For the target word “orange”, there are two strong prototypes, namely “fruit” and “color”. Depending on the probe, they are either tied, or “fruit” is the most prototypical item, with “color” being the second. This aligns with data by Battig and Montague [3] and Overschelde et al [46], where more than 80% of subjects named both “fruit” and “color” as hypernyms of “orange”.
We now address the question whether all relations show a prototypicality effect. We first consider the responses with the strongest prototypicality, namely, the zero response entropy, where the same single word was the only response of all participants.
Table 3: Responses with zero response entropy.
Relation Ratio of responses with $\mathcal{R}$=0 Length
Antonymy (ANT) 6.11% (50/819) 4.28
Synonymy (SYN) 2.44% (36/1477) 5.25
Holonymy (HOL) 1.31% (17/1302) 5.93
Hypernymy (HYP) 1.14% (55/4844) 5.59
Hyponymy (HPO) 0.16% (2/1240) 7.46
Meronymy (MER) 0.12% (1/864) 7.62
Table 3 lists the ratio of such responses, along with the average number of words in responses. According to this metric, antonymy shows by far the strongest prototypicality at 6.11% of all responses, followed by synonymy at 2.44%. We can observe that meronymy and hyponymy almost never show responses with the strongest prototypicality (only once for meronymy and twice for hyponymy). This might be related to the fact that these two relations happen to also have more relata in the human responses (more than seven words on average) than other relations, which have an average of around five.
Figure 2 shows the distributions of response entropies $\mathcal{R}$ across relations [^15] Distributions are significantly different from each other for all relation pairs except for holonymy-hypernymy and holonymy-synonymy relation pairs, as established by Mann-Whitney U tests with $\alpha$=0.05.. For visibility reasons, we excluded zero scores from Figure 2161616However, zero scores were not excluded when applying the statistical tests..
Figure 2: Response Entropy (excluding zero cases). Boxes enclose second and third quantiles, with the mean shown as orange lines. A lower score shows a stronger prototypicality.
We can see that the strongest prototypicality effect by far is again observed for antonymy, at a mean of 0.83. Synonomy shows the second strongest prototypicality with a mean of 0.90, with hypernymy and holonymy somewhat less prototypical. Hyponymy and meronymy are again at the other extreme, with $\mathcal{R}$ means above 0.95, close to the maximum of 1, and standard deviations lower than those of other relations. This means that most distributions for these two relations are close to uniform. Combining these observations, we conclude that antonymy shows a strong prototypicality effect, but there is hardly any prototypicality effect for hyponymy and meronymy. For our planned prototypicality evaluation of model, we will not use hyponymy and meronymy relations, but only hypernymy, holonymy, antonymy, and synonymy.
If a human response shows a $\mathcal{R}$ of 1, the probe is unable to elicit any prototypical response. We also remove probes whose responses show a $\mathcal{R}$ of 1.0 from these four relations in the evaluation of prototypicality. We find 401 such cases for hypernymy, 199 for hyponymy, 93 for synonymy, 79 for holonymy, 65 for meronymy, and 15 for antonymy. In addition, words that are out of model’s vocabularies will introduce an atrificial deflation in evaluation. We therefore further discard responses that include any word that is not in the intersection vocabulary of the models tested. This results in a total of 4,331 responses which we can use in our prototypicality experiments: 2,406 for hypernymy, 805 for holonymy, 448 for antonymy, and 672 for synonymy.
7 Model Experiment
For soundness, completeness, and distinguishability, we use all 10,546 probes from Table 1. For model prototypicality, we use 4,331 probes, as described above. For symmetry and asymmetry, we use 17,608 probes. The number is higher than 10,546 because of the trick probes we added (cf. Section 5.2). For symmetry and asymmetry, the values of $k$ we use are 1, 5, and 10.
7.1 Target Models
We use BERT as one of the MLMs because it is widely used in previous work (we use the cased version). In addition to BERT, we also chose RoBERTa [33] and ALBERT [29], as this allows us to quantify the effect of training objectives and architectures.
In this paper, we also present experiments with CLMs. The most important condition for comparability between neural models is that their vocabularies are as similar as possible. To select the most suitable model, we experimentally determine the Jaccard similarity between two models’ vocabulary. The average Jaccard similarity within the set of MLMs (BERT-RoBERTa, BERT-ALBERT, ALBERT-RoBERTa) is 0.40. We then compare the similarity between each of the two CLMs (OPT and Llama-2) with the three MLMs. When we consider the similarity of OPT with the three MLMs, the average Jaccard similarity (within the new set of four models) increases from 0.40 to 0.50171717Part of the reason for this increase is the fact that OPT and RoBERTa share vocabularies.. For Llama-2, the pairwise similarity drops from 0.40 to 0.32. Based on these results, we choose OPT over Llama-2.
In order to study how model size impacts the learning of semantic relations, we also use models of different sizes within each model family. For BERT and RoBERTa, the two variants we use are of similar sizes, but for OPT, the range of sizes we experiment with is much larger, namely three orders of magnitude. For ALBERT, the size range is somewhere in the middle.
Table 4 lists the statistics of the target models we use.
Table 4: Statistics of our target models.
Abbr. Model #Parameters Vocabulary Size Pretraining Set Size
B1 BERT-base 110M 28,996 16GB
B2 BERT-large 340M 28,996 16GB
A1 ALBERT-base 12M 30,000 16GB
A2 ALBERT-large 18M 30,000 16GB
A3 ALBERT-xlarge 60M 30,000 16GB
A4 ALBERT-xxlarge 235M 30,000 16GB
R1 RoBERTa-base 125M 50,265 160GB
R2 RoBERTa-large 355M 50,265 160GB
O1 OPT-125M 125M 50,265 800GB
O2 OPT-350M 350M 50,265 800GB
O3 OPT-1.3B 1.3B 50,265 800GB
O4 OPT-2.7B 2.7B 50,265 800GB
O5 OPT-6.7B 6.7B 50,265 800GB
O6 OPT-13B 13B 50,265 800GB
O7 OPT-30B 30B 50,265 800GB
O8 OPT-66B 66B 50,265 800GB
7.2 Dealing with Determiner Bias
Some probes require indefinite determiners before the nouns that are to be predicted. Keeping determiners fixed in these probes would introduce a bias towards words with an initial vowel or consonant. For each probe that requires indefinite determiners, we therefore probe the models twice, once with “an” inserted into the probe and the other time with “a” inserted into the probe. From the responses, we create a new distribution over the vocabulary, which is the weighted sum of the two distributions yielded by the two probes. The weights correspond to the relative frequencies of “an” and “a” in the Corpus of Contemporary American English [13]. This synthetic distribution can then be treated as the tested model’s prediction to the probe.
7.3 Statistical Test
Throughout this study, statistical differences in soundness, symmetry, and asymmetry metrics have been tested using McNemar’s test with $\alpha=0.05$, as these metrics are binary for each target word or each tuple. For completeness and prototypicality, the Wilcoxon signed rank test is used, with $\alpha$=0.05. For distinguishability, no known test exists so we do not test for significance.
8 Results and Analyses
Before presenting the results of each metric, we examine the general characteristics of responses from each agent. Models produce more OOR responses than humans. The best model, RoBERTa-large (R2), returns the first non-OOR word on average at rank 40.6; recall that this number was 1.5 for humans. Even for the best model, 64% of response lists consist solely of OOR words (all other models’ numbers are even higher). The number for humans was 20%. However, some of the non-OOR responses we received both from models and from humans might be due to holes in our gold standard.
Disregarding OOR responses, Figure 3 shows the distribution of the relation between the first non-OOR word returned by agents and the target word, compared to the gold standard (the prompted-for relation).
Figure 3: The Distribution of the relations that the first non-OOR word is in with the target word for each model family, disregarding all OOR responses.
Humans’ distribution closely resembles the gold standard distribution. However, they slightly underestimate the proportions of holonymy and meronymy. Models’ distributions are less similar to the gold standard, with a clear tendency to underestimate the proportion of hypernymy and meronymy and overestimate the proportion of antonymy. The distribution of relations that is most different from the gold distribution comes from OPT, which performs far worse than the three MLMs. We now present a more detailed analysis as made possible by our specialized metrics.
8.1 Soundness
Figure 4 shows the result of the soundness evaluation for our models, along with the human ceiling.
Figure 4: Results for soundness.
In the graphs, we show significance using dotted lines. A dotted line indicates that the test established statistically significant differences between all models above and below the line [^18] Please note that this is not equivalent to saying that all models between neighbouring dotted lines are statistically indistinguishable. This may or may not be the case for any pair; the notation we use is a simplification in that it cannot express this aspect..
Humans (H) show high performance on antonymy ($\mathcal{S}=0.90$), with lower scores for the other relations ($0.750.63<\mathcal{S}<$ 0.75, with an average of 0.66). The performance of all models remains far below that of humans: for most relations, even the best model score is less than half the human score.
As for individual models, we can see that for all relations RoBERTa-large (R2) is either the significantly best model, or in the best-performing group. The one relation where models perform relatively well is antonymy, where the best models achieve $\mathcal{S}>0.45$, whereas in other relations best values typically lie around $\mathcal{S}=0.25$. CLMs overall perform less well, with the best scores achieved by any CLM ranging from $\mathcal{S}=0.04$ for synonymy to $\mathcal{S}=0.38$ for antonymy.
8.2 Completeness
Figure 5 shows the results for completeness.
Figure 5: Results for completeness.
Again, we observe a large gap between models and humans, as was the case for soundness earlier. For relations except for antonymy, the human $\mathcal{C}$ scores range from 0.38 to 0.49 and are therefore overall much lower than the human $\mathcal{S}$ scores, where even the lowest score was above 0.50.
$\mathcal{C}$ scores for models remains below 0.25 for the five relations except antonymy. All other trends are similar to those for soundness: antonymy stands out again as a relation with high completeness, for both humans and models. The overall best performer is again RoBERTa-large (R2).
Among all relations evaluated, antonymy stands out. Both $\mathcal{C}$ and $\mathcal{S}$ scores of antonymy are higher than other relations, for all agents evaluated.
We conclude from the results for OOR-rate, soundness, and completeness that the models only acquire a limited ability to perform relata prediction, which is far below the human ceiling.
8.3 Symmetry and Asymmetry
Let us consider the two symmetric relations first. For $k$=5, Figure 6(a) shows the evaluation results for $\mathcal{M}$ scores [^19] Results for $k$=1 and $k$=10 are given in Appendix B. Overall, we observed similar trends for all values of $k$..
(a) Symmetry.
(b) Asymmetry.
Figure 6: Results for symmetry and asymmetry.
The humans achieve $\mathcal{M}$=0.88 for antonymy and $\mathcal{M}$=0.65 for synonymy. The models similarly perform better for antonymy than they do for synonymy. For antonymy, the best-perfoming group includes BERT-large (B2; $\mathcal{M}=0.62$) and RoBERTa-large (R2; $\mathcal{M}=0.6$)). For synonymy, it includes ALBERT-xxlarge (A4; $\mathcal{M}=0.30$), BERT-large (B2; $\mathcal{M}=0.28$) and RoBERTa-large (R2; $\mathcal{M}=0.27$). Again, for both symmetric relations, the best CLM is left far behind.
We now move to the results of asymmetric relations. Figure 6(b) shows the evaluation results for $\mathcal{A}$ scores, again for $k$=5. The models generally perform better on hypernymy than they do on hyponymy. They also perform better on holonymy than on meronymy. Across all relations, RoBERTa-large (R2) consistently ranks as the top model or is in the best-performing group. However, no results above $\mathcal{A}$=0.40 were measured.
In conclusion, all models tested recognized the asymmetry of four asymmetric relations, and the symmetry of synonymy, only to a limited extent. In contrast, the best set of models recognized the symmetry of antonymy to a relatively high extent, approaching human performance.
8.4 Prototypicality
Figure 7 gives the evaluation results for $\mathcal{P}$ scores. Remember that prototypicality is not reported for hyponymy and meronymy, as we have established in Section 4 that the human responses showed no prototypicality effects for these relations.
Figure 7: Results for prototypicality.
All models for hypernymy, holonymy, and synonymy achieved results below $\mathcal{P}$=0.3. The best results ($\mathcal{P}$=0.26 for holonymy, $\mathcal{P}$=0.23 for synonymy and $\mathcal{P}$=0.21 for hypernymy), were obtained by RoBERTa-large (R2). Antonymy again is the positive outlier relation, with best results ranging around $\mathcal{P}$=0.40 (achieved by ALBERT-xxlarge (A4), RoBERTa-base (R1), BERT-large (B2) and RoBERTa-large (R2)). Therefore, we can conclude that the models only learn a limited degree of prototypicality for hypernymy, holonymy, and synonymy, while they perform relatively well for antonymy.
8.5 Distinguishability
Figure 8: Distinguishability curves (above) and AuDC (below).
Figure 8 shows the distinguishability curves for humans and models, with corresponding AuDC values in the lower part of the figure. We can see in the human results that as $p$ increases, the theoretical maximum of $\eta(p;m)$=30 is kept up until after $p=0.5$. The curve then gradually descends, and reaches zero around $p=0.9$. The curves for the models show a faster descent than that for humans; zero $\eta(p;m)$ is reached around $p=0.65$. This means that the upper bound for models only slightly exceeds the lower bound for humans, suggesting a substantial difference between models and humans in their ability to distinguish relation pairs.
Figure 9: Distinguishability matrices of MLM, CLM, and humans.
Now let us move on to the lower part of Figure 8, which shows the AuDC values. The AuDC tells us how many relation pairs out of 30 are distinguished by an agent, on average. For humans, the $\mathrm{AuDC}$ value is 21.3, whereas models are only able to distinguish 4.0 to 9.8 pairs, fewer than a third of all relations. We can therefore safely say that the meta-relational knowledge of all models we tested is unsatisfactory.
Plotting all pairwise $\mathcal{D}$ scores for an agent, we can create the agent’s distinguishability matrix. This serves to examine more deeply which relations are mistaken for which other ones. The relation given at the row position is the prompted relation, and the relation given at the column position is the relation the model responded with. The depth of grey shade expresses the degree of distinguishability, lighter cells indicating that the row relation is more often confused with the column relation.
Figure 9 presents $\mathcal{D}$ scores for entire model families, calculated as averages over all model variations of MLM (left) and CLM (middle), in comparison to the human ceiling (right) [^20] The complete list of distinguishability matrices for all model variants is provided in Appendix C.. Humans, despite their generally good ability to distinguish semantic relations, tend to confuse hyponymy and hypernymy with synonymy, as can be seen from the lighter cells at the leftmost bottom ((SYN, HYP) and (SYN, HPO)) and at the rightmost top ((HYP, SYN) and (HPO, SYN)). This agrees with findings by Chaffin and Clark: humans perceive synonymy as being close to hypernymy and hyponymy (cf. Section 2). We also observe that humans’ $\mathcal{D}$ scores for hypernymy versus both holonymy and meronymy (lighter cells, (HOL, HYP) and (MER, HYP) in the leftmost middle) are relatively low as well. This aligns with the theoretical assumption by Cruse [12], Winston et al [64] and Joosten [27] that there are similarities between holonymy and hypernymy.
What stands out in both the models’ and humans’ distinguishability matrices are the high scores when antonymy is the relation prompted (visible in the antonymy row). However, in the antonymy column, which shows the cases where the model responded with antonyms although the prompt was associated with a non-antonymy relation, high $\mathcal{D}$ scores are observed only for humans, but not for models.
To sum up, we observe that models can distinguish antonymy from other relations, but not other relations from antonymy. Antonymy is also the relation for which the models show the highest scores in all metrics we introduced: not only for distinguishability, but also for soundness, completeness, symmetry, and prototypicality. These observations lead us to conclude that there must be some kind of antonymy bias: irrespective of the relation prompted, models consistently prefer antonyms of the target words. We now show an example of this antonymy bias.
Table 5: Top three words in the prediction of models given the probe “an answer is similar to [det] [v]. Antonyms are shown in boldface.
Abbr. Agent Top 1 Top 2 Top 3
H Humans response reply solution
B1 BERT-base question answer statement
B2 BERT-large question statement answer
R1 RoBERTa-base question yes answer
R2 RoBERTa-large question answer argument
A1 ALBERT-base question query error
A2 ALBERT-large question answer query
A3 ALBERT-xlarge question inquiry query
A4 ALBERT-xxlarge question reply query
[…]
Humans correctly produced only synonyms for this probe, but all models except for the 4 smallest OPT models return “question” as the first response; O1, O3 and O4 return “question” as their first non-OOR response, while OPT-350M (O2) returns only non-OOR words in the top three. This suggests that the relation of “question” and “answer” is learned firmly by the models, particularly the MLMs.
8.6 Model Size Analysis
Table 6: Performance difference between the largest and the smallest model for four model families, across relations. The largest
Metric Avg. Type BERT RoBERTa ALBERT OPT
Soundness ($\Delta\mathcal{S}$) Micro +.$+.04$ +.$+.09$ +.$+.11$ +.$+.07$
Macro +.$+.06$ +.$+.07$ +.$+.14$ +.$+.08$
Completeness ($\Delta\mathcal{C}$) Micro +.$+.05$ +.$+.04$ +.$+.09$ +.$+.06$
Macro +.$+.06$ +.$+.04$ +.$+.12$ +.$+.08$
[As/S]ymmetry ($\Delta\mathcal{A}\mid\Delta\mathcal{M}$) Micro +.$+.05$ +.$+.06$ +.$+.12$ +.$+.10$
Macro +.$+.07$ +.$+.07$ +.$+.16$ +.$+.12$
Prototypicality ($\Delta\mathcal{S}$) Micro +.$+.05$ +.$+.07$ +.$+.09$ +.$+.06$
Macro +.$+.06$ +.$+.07$ +.$+.14$ +.$+.08$
Distinguishability ($\Delta$AuDC) n.a. $+1.23$ $+1.42$ $+4.36$ $+4.59$
According to the scaling law, models with large sizes should outperform their smaller counterparts, so one should see only positive differences when model size increases. We examine whether this holds for our tasks. Table 6 presents the differences in performance between the largest model and the smallest model in each model family, for all metrics we consider [^22] The term “[As/S]ymmetry” refers to symmetry scores for symmetric relations and asymmetry scores for asymmetric relations.. Except for AuDC, all metrics are reported as two different averages. Macro differences are averaged over relations, whereas micro differences are averaged over individual scores per target word.
As all differences in all metrics shown in Table 6 are positive, the largest model outperforms the smallest model in the same family in every case. If there is a performance boost, however, its size cannot be predicted from the model size increase alone. The same increase in size (330 million) generally affords RoBERTa a higher degree of improvement than it does BERT. The comparison between OPT and ALBERT is also illustrative in this respect. Despite ALBERT’s smaller increase in model size (from 12 to 235 million) when compared to that of OPT (from 125 million to 65 billion), ALBERT always experiences a larger performance boost than OPT. When the performance differences are broken down into individual relations (results not shown here, but in Appendix D), we find that for most model families, antonymy is the relation that shows the highest improvement among all relations, ranging from 0.14 to 0.34. In contrast, the highest score reached for any other relation is only 0.16.
So far, we have only looked at comparisons of largest and smallest model in a family. We now look at smaller increases from one model to the next-larger model within its model family. We count the number of cases when a smaller model performs better than its next-larger model for all metrics and all relations [^23] If a model family has $n$ members, we perform $n$-1 tests. A table of detailed results can be found in Appendix D.. If this happens, the function between model size and performance is non-monotonic.
OPT is the family with the most non-monotonic behaviour out of the model families we consider: smaller models outperform the larger at least once for almost every relation and metric pair (except for hyponymy soundness, where we found no significance). Non-monotonic size-performance behaviour can also be observed for ALBERT in the following cases: all metrics for hypernymy, hyponymy completeness and asymemtry, holonymy soundness and prototypicality, and antonymy symmetry.
To sum up, large models generally outperform small models. However, the models’ performance does not always increase monotonically with model size. Therefore, for the learning of semantic relations, model size is not all the matters.
8.7 Pretraining Task Analysis
In factual probing tasks, MLMs has been shown to outperform CLMs [49, 8, 42], as we have discussed in Section 2. We now verify whether this phenomenon also holds for semantic relations.
We first calculate differences between pairs of best-performing MLMs and best-performing CLMs across model families, per metric. The results are presented in Table 7.
Table 7: Difference between the best models pretrained on different tasks (the best MLM minus the best CLM). All within-metric differences are statistically significant.
Relation Soundness $\Delta\mathcal{S}$ Completeness $\Delta\mathcal{C}$ [As/S]ymmetry $\Delta\mathcal{A}\mid\Delta\mathcal{M}$ Prototypicality $\Delta\mathcal{S}$
HYP $+$.15 $+$.06 $+$.07 $+$.09
HPO $+$.20 $+$.09 $+$.09 n.a.
HOL $+$.12 $+$.07 $+$.05 $+$.11
MER $+$.18 $+$.13 $+$.14 n.a.
ANT $+$.10 $+$.10 $+$.10 $+$.07
SYN $+$.17 $+$.14 $+$.24 $+$.19
As there are only positive performance differences in Table 7, it is never the case that the best CLM outperforms the best MLM. The best-performing MLM is always either RoBERTa-large (R2) or BERT-large (B2), whereas the best CLM is always either OPT-13B (O6), OPT-30B (O7), or OPT-66B (O8); different conditions produce different pairs. For AuDC, we find that the best-performing pair is established by OPT-66B and RoBERTa-large, with a numerical difference of 1.02 in favour of the MLM.
We want to point out that in all cases, the losing CLM model is larger than the winning MLM model. The smallest size difference observed between the pair of best MLM and best CLM (O6 and R2) is 12 billion parameters.
We also observe that the performance difference between MLM and CLM varies according to relations. The largest difference can be observed for synonymy, followed by meronymy, hyponymy, hypernymy, antonymy and holonym, in this order. This is exactly the opposite observation that we saw for model size, where antonymy profited most from larger sizes.
To gather further evidence, we run every MLM model against every CLM model, and count the number of pairs where the MLM is significantly better. Table 8 shows the results.
Table 8: Number of MLM-CLM pairs where the MLM significantly outperforms the CLM, out of a total of 64.
Relation Soundness Completeness. [As/S]ymmetry Prototypicality
HYP 48 46 40 48
HPO 57 60 60 n.a.
HOL 44 41 29 41
MER 61 57 56 n.a.
ANT 49 47 47 50
SYN 61 62 63 62
The performance of the MLM is significantly better for at least 41 out of 64 MLM-CLM pairs, except for holonymy under asymmetry, where it is 29 pairs. For AuDC, as there is no statistical test, we report MLM-CLM pairs that yield numerical differences: there are 40 such pairs. Overall, this confirms the general superiority of MLMs over CLMs, particularly if we consider that in these pairs there are many small MLMs outperforming larger CLMs. The gap between MLMs and CLMs cannot be bridged by increases in model sizes, despite the tendency of larger models to learn semantic relations better.
8.8 Word Frequency Analysis
BERT is known to achieve higher accuracy scores in hypernym prediction tasks when the target word is frequent [50]; it is therefore prudent to perform a correlation analysis of results and word frequency. If the frequencies of target words and relata are partially responsible for the performance of a model, there should be a positive correlation between performance and frequency.
For every model, we use the rank correlation metric Spearman’s $\rho$ to compare word-frequency metrics that we derive independently from COCA, against all metrics except AuDC. We exclude AuDC from this analysis, as AuDC is a metric that is not lexically determined. As soundness, completeness, and prototypicality are summary statistics calculated across target words, we first need to recompile individual scores per target word. These scores can be obtained using Equations (11), (13) and (22). We then correlate the scores of each target word in these three metrics with the target word’s frequency.
Soundness and completeness are relations that involve relata sets. We therefore also need to consider the frequency of relata, not only of target words [^24] Prototypicality scores do not use relatum sets as gold standard. Thus, we only report correlation with target words for prototypicality.. We calculate the correlation between the scores against the average and maximum frequency of relata in each gold relatum set. We choose the average and maximum because we need to consider the relatum set as a whole, as this is how the sets are used in the calculation of these metrics.
Asymmetry and symmetry require special treatment because they are defined on tuples involving two words (the target word and the relatum). Agents might be unfairly advantaged in recognizing symmetry or asymmetry by word frequency effects in two cases: 1) if both words in the tuple are common and 2) if one word in the tuple is far more common than the other. We use two metrics: average frequency, which can guard against the first case, and absolute frequency difference, which can guard against the second. We first recalculate the symmetry and asymmetry scores per tuple, using Equations (15) and (17), and then correlate them with the average frequency and the absolute frequency difference.
This allows us to determine correlation for certain metric and relation combinations: For soundness and completeness, there are coefficients for all six relations. For symmetry, there are coefficients for the two symmetric relations. For asymmetry, there are coefficients for the four asymmetric relations. For prototypicality, there are coefficients for hypernymy, holonymy, antonymy, and synonymy. Consequently, the total number of coefficients is $16\times(2\times 6+1\times 2+1\times 4+1\times 4)=352$. The results are that correlations for all models have medians below 0.30 across all metrics and relations considered [^25] Details can be found in Appendix E.. Hinkle et al [25] regard correlations below 0.30 as negligible. We only conclude that there is a certain influence of word frequencies on the performance of semantic relation tasks, as was to be expected, but we believe that it is unlikely to be the defining factor in semantic relation learning.
9 Limitations
Our work has some methodological limitations. We adopt prompt-based probing as our core method, but all prompt-based methods suffer from a high dependency on specific prompt design [50, 15, 7]. We counteract this dependency by using several different prompts for each semantic relation, but we cannot be sure that this is enough. Cao et al [8] presents a method for the mitigation of prompt dependency in evaluations, which we implemented and applied to our results, but which resulted in little difference [^26] All results reported in the paper were therefore given in their original form, i.e., without mitigation..
A possible way to investigate probe dependency more thoroughly is to determine whether the scores from different prompts are heteroscedastic or homoscedastic. Heteroscedasticity is the property of several samples to have a different variance [5]. Preliminary results of this analysis are given in Appendix F. We found that for all models, some prompts present heteroscedasticity under certain metrics, but for humans, the evidence supporting heteroscedasticity is insufficient.
In the general case, prompt dependency remains an unsolved question. We suspect that some models may use linguistic expressions in certain prompts as shortcuts when solving semantic relation task. For improving the evaluation methodology presented here, the identification of such shortcut expressions is the next step for the mitigation of prompt dependency, which should secure more stable results.
In our word frequency analysis above, we found little correlation between scores and word frequency, but we only examined unigram frequency of individual words. Unigram frequency cannot use any information of syntagmatic or paradigmatic relationships between two words. Using frequency-based metrics that are able to exploit such information may lead to a different conclusion. Particularly, more complex frequency measures that capture syntagmatic or paradigmatic relationships would be desirable. However, they require a precise definition of such measures and carefully controlled experiments, which are aspects beyond the scope of the present study.
Our evaluation also suffers from the fact that we only prompt with individual words and disregard subwords. The vocabulary of PLMs is a mixture of words and subwords, where frequent words remain as they are and less frequent words are split into subwords. Recent models like Llama [60] minimize their vocabulary size, instead including more subwords in their vocabulary. This may disadvantage them in our evaluation (and we even excluded Llama based on low vocabularly overlap). As a result, we were unable to gain knowledge about other CLMs, some of which are more widely used than OPT.
10 Conclusion
Current PLMs are commonly used for a wide range of tasks, so it is important to gain a comprehensive understanding of their linguistic abilities. This study focuses on the semantic relation knowledge of MLMs and CLMs. In particular, it explicates the gap in semantic relation knowledge between current PLMs and humans. Our contributions are as follows.
  1. 1.
    We presented a prompt-based probing evaluation methodology that covers six aspects of semantic relation knowledge, namely soundness, completeness, symmetry, asymmetry, prototypicality, and distinguishability. Two of these metrics are novel, namely those for prototypicality and for distinguishability. We also employed established metrics in a new context.
  2. 2.
    Using these evaluation methods, we conducted a comprehensive evaluation of the above-mentioned aspects for six semantic relations, five of which were never empirically tested in probing experiments before.
  3. 3.
    We established the first human gold standard for prototypicality. For the other aspects, where we constructed gold standards from existing lexicographic data, we established the first human ceiling, which can be used in comparisons with automatic models.
Our experiment afforded far more conditions and distinctions than previous studies, including a human ceiling and a comparison of CLMs with MLMs. In this way, we arrive at a richer characterization of PLMs’ capabilities with semantic relations than was possible before. Our main result is that PLMs fall short of achieving human-level performance on the extensive semantic relation tasks defined here.
We experimentally studied the prototypicality effect as displayed by humans for the six relations. We were able to confirm prototype effects for hypernymy and antonymy that have been experimentally studied in the literature. We also found a prototype effect for holonymy and synonymy that was not known before. One of our most imporantant findings with respect to prototypicality was that hyponymy and meronymy showed little prototypicality.
We also found that when it comes to learning semantic relations, a large model size does not guarantee better performance in our tasks. The type of PLMs also matters. MLMs consistently outperformed CLMs across all metrics for all semantic relations. It therefore seems likely that the bidirectional context utilized by MLMs is a crucial factor in learning semantic relations.
Out of all relations, antonymy is the one where both humans and models performed best, a result which is stable across all metrics. We also observed an antonymy bias operating in all models: while they were able to distinguish antonymy from non-antonymy, they often misrecognize non-antonymy as antonymy. Humans performed well with antonymy in both directions, as has been anticipated in previous work [12, 27, 9]. Since the antonymy bias appears across models, it may be attributed to some distributional characteristics of the antonymy relation which make it fundamentally different from the other relations. For example, although antonymy is mainly a paradigmatic relation, it also has a strong syntagmatic aspect: antonymy pairs commonly co-occur in conjunction structures such as “ascent and descent” and “either day or night”. While the other relations may also display a mixture of paradigmatic and syntagmatic features, the effect is certainly not as strong as for antonymy. In future work, it might be fruitful to study how the dual distributional nature of antonymy affects models’ recognition of it. Such insights might lead to better learning methods for the other relations.
Our study aims to contribute towards a future where PLMs can better understand language. The fact that PLMs struggle to understand several aspects of semantic relations contradicts the superiority of PLMs as observed in many NLP tasks. This superiority is commonly attributed to the assumption that PLMs are able to efficiently encode general semantic and linguistic knowledge. Our work, which is a thorough investigation of this ability, showed that this is evidently not so. Some other explanation for the good performance should be sought. Our methodology is able to substantiate such doubts; in general, it captures effects not seen before. We therefore consider it a prism through which to see the truth more clearly.
Appendix A All Prompts
Table 9: All prompts used in this research. Presented per relation.
Relation Prompt
HYP (7) [det] [w] is a type of [det] [v]
[det] [w] is a kind of [det] [v]
the word [w] has a more specific meaning than the word [v]
[det] [w] is [det] [v]
[det] [w] is a specific case of [det] [v]
[det] [w] is a subordinate type of [det] [v]
the word [w] has a more specific sense than the word [v]
HPO (4) my favorite [w] is [det] [v]
[det] W, such as [det] [v]
the word [w] has a more general meaning than the word [v]
the word [w] has a more general sense than the word [v]
HOL (7) [det] [w] is a component of [det] [v]
[det] [w] is a part of [det] [v]
[det] [w] is contained in [det] [v]
[det] [w] belongs to constituents of [det] [v]
[det] [w] belongs to parts of [det] [v]
[det] [w] belongs to components of [det] [v]
[det] [w] is a constituent of [det] [v]
MER (6) constituents of [det] [w] include [det] [v]
components of [det] [w] include [det] [v]
parts of [det] [w] include [det] [v]
[det] [w] consists of [det] [v]
[det] [w] has [det] [v]
[det] [w] contains [det] [v]
ANT (9) it is not likely to be both [det] [w] and [det] [v]
[det] [w] is the opposite of [det] [v]
the word [w] has an opposite sense of the word [v]
it is impossible to be both [det] [w] and [det] [v]
the word [w] has a meaning that negates the meaning of the word [v]
it is [det] [w] so it is not [det] [v]
the word [w] has an opposite meaning of the word [v]
if something is [det] W, then it can not also be [det] [v]
the word [w] has a sense that negates the sense of the word [v]
SYN (7) [det] [w] is also known as [det] [v]
[det] [w] is often referred to as [det] [v]
the word [w] has a similar meaning as the word [v]
[det] [w] is similar to [det] [v]
the word [w] means nearly the same as the word [v]
[det] [w] is indistinguishable from [det] [v]
[det] [w] is also called [det] [v]
Appendix B Results of Symmetry and Asymmetry
(a) Symmetry.
(b) Asymmetry.
Figure 10: Results for symmetry and asymmetry when $k=1$.
(a) Symmetry.
(b) Asymmetry.
Figure 11: Results for symmetry and asymmetry when $k=10$.
Appendix C Confusion Matrices
Figure 12: Distinguishability matrices of all MLMs
Figure 13: Distinguishability matrices of all CLMs
Appendix D Model Size Difference per Relation and Metric
Table 10 shows the performance difference between the largest and the smallest model for four model families. Table 11 shows the cases where a model significantly outperforms its next-larger counterpart.
Table 10: Performance difference between the largest and the smallest model for four model families. All differences in metrics where tests can be performed are significant unless they appear in a bracket. The maximum within model families and metrics is boldfaced.
Metric Relation BERT RoBERTa ALBERT OPT
Soundness HYP .02 .12 .07 .06
HPO (.00) .11 .11 .09
HOL .04 (.02) .08 .07
MER .03 .05 .12 (.01)
ANT .19 .04 .33 .25
SYN .05 .10 .14 .02
Completeness HYP .03 .04 .06 .06
HPO .05 .02 .06 .06
HOL .04 .04 .07 .06
MER .03 .04 .11 .02
ANT .16 .04 .30 .24
SYN .06 .07 .12 .03
[As/S]ymmetry HYP .04 .07 .11 .12
HPO (.00) .02 .02 .03
HOL .05 .06 .16 .10
MER .05 .05 .16 .06
ANT .15 .09 .29 .34
SYN .11 .13 .22 .04
Prototypicality HYP .03 .08 .06 .06
HOL .05 .07 .06 .05
ANT .14 (.03) .25 .22
SYN .05 .08 .12 .02
Distinguishability n.a. 1.23 1.42 4.36 4.59
Table 11: Number of smaller models who significantly outperform the next-largest model.
Metric Model Family (# Models) HYP HPO HOL MER ANT SYN
Soundness BERT (2) 0 0 0 0 0 0
RoBERTa (2) 0 0 0 0 0 0
ALBERT (4) 1 0 1 0 0 0
OPT (8) 3 0 2 1 1 1
Completeness BERT (2) 0 0 0 0 0 0
RoBERTa (2) 0 0 0 0 0 0
ALBERT (4) 1 1 0 0 0 0
OPT (8) 3 2 2 1 1 1
[As/S]ymmetry BERT (2) 0 0 0 0 0 0
RoBERTa (2) 0 0 0 0 0 0
ALBERT (4) 1 1 0 0 1 0
OPT (8) 3 1 1 1 1 1
Prototypicality BERT (2) 0 0 0 0 0 0
RoBERTa (2) 0 0 0 0 0 0
ALBERT (4) 1 n.a. 1 n.a. 0 0
OPT (8) 3 n.a. 2 n.a. 1 2
Appendix E Word Frequency Correlation
Figure 14 presents the results of all coefficients per metric and relation.
Figure 14: Spearman’s $\rho$ between soundness, completeness, asymmetry, symmetry, and prototypicality against word-frequency metrics.
Appendix F Performance Heteroscedasticity Introduced by Prompts
Prompts are known to influence in spectrum of tasks [15, 7]. We wonder if this finding holds in semantic relation tasks as well. In order to figure it out, we assess whether prompts introduce performance heteroscedasticity, where performance variances differ significantly across prompts, for agents. The heteroscedasticity indicates that an agent’s performance originates from populations with varying variances when using different prompts. Therefore, observing the heteroscedasticity suggests the influences of prompts for the agent.
We use Levene’s test to determine heteroscedasticity, interpreting the results at a significance level of 0.05. We target soundness, completeness, and symmetry. Prototypicality is excluded because not all prompts are used in its calculation (c.f. Section 6.2). For each relation and its prompts, we calculate metrics as if each prompt is the only one in the set. Thus, for a given metric and relation, we obtain $N$ sets of results with $N$ prompts. We apply Levene’s test on these $N$ sets.
Results are presented in Table 12. Only models exhibit prompt heteroscedasticity, with the number of such models varying by metrics and relations. No sufficiently strong evidence is found supporting the heteroscedasticity in human performance for any metrics and relations.
Table 12: Agents, given per metric and relation, present performance heteroscedasticity introduced by prompts. The column “TOTAL” shows the number of such agents out of 17 agents (16 models + humans).
Metric Relation Agents TOTAL
Soundness HYP All models 16
HPO A1, A2, A3, A4, B1, B2, R1, R2, O2, O4, O5, O6, O7, O8 14
HOL All models 16
MER A3, B2, R1, R2, O2, O3, O6 7
ANT A1, A2, A3, A4, B1, B2, O1, O2, O3, O4, O5, O6, O7, O8 14
SYN A2, A4, B2, R2, O1, O4, O5, O6, O7, O8 10
Completeness HYP All models 16
HPO A1, A2, A4, B1, B2, R2, O1, O2, O3, O4, O5, O6, O7, O8 14
HOL All models 16
MER A3, R1, R2, O2, O3, O6, O7 7
ANT A1, A2, A3, A4, B1, B2, O1, O2, O3, O4, O5, O6, O7, O8 14
SYN A4, B1, R2, O1, O3, O4, O5, O6, O7, O8 10
[As/S]ymmetry HYP All models 16
HPO All models 16
HOL ALL model 16
MER A3, B2, R1, R2, O1, O2, O3, O4, O6, O7, O8 11
ANT A1, A3, A4, O1, O2, O3, O4, O5, O6, O7, O8 11
SYN A1, A2, A4, B1, B2, R1, R2, O1, O3, O4, O5, O6, O7, O8 14